# On the semi-additivity of the $1/2$-symmetric caloric capacity in the plane

### v3

@@ Abstract.
In this paper we study properties of a variant of the $1/2$-caloric capacity, called $1/2$-symmetric caloric capacity. The latter is associated simultaneously with the $1/2$-fractional heat equation and its conjugate. We establish its semi-additivity in $\mathbb{R}^{2}$ and, moreover, we compute explicitly the $1/2$-symmetric caloric capacity of rectangles, which illustrates its anisotropic behaviour.
@@ 1. Introduction
The study of removable subsets for caloric equations done by Mateu, Prat & Tolsa in [MPrTo] and by Mateu & Prat in [MPr] introduces, naturally, caloric capacities associated to such PDE’s. In the first reference $(1,1/2)$-Lipschitz caloric capacity was introduced, as well as the notion of equivalence between the nullity of this quantity and the removability of compact subsets for the heat equation, i.e. the one associated with the differential operator $\Theta:=(-\Delta_{x})+\partial_{t}$, where $(x,t)\in\mathbb{R}^{n}\times\mathbb{R}$. In [MPr], Mateu & Prat studied the corresponding caloric capacities associated with the fractional heat equation. That is, the equation associated with the pseudo-differential operator $\Theta^{s}:=(-\Delta_{x})^{s}+\partial_{t}$, for $10<s<$ 1. The authors distinguished the cases $s=1/2$, $11/2<s<$ 1 and $20<s<1/$ 2, focusing mainly on the first. The study for $s=1/2$ was quite successful, obtaining a removability result for the $\Theta^{1/2}$-equation, as it was done for the heat equation. For instance, if $f$ is a solution of the $1/2$-heat equation in $\mathbb{R}^{n+1}\setminus{E}$ satisfying
$\|f\|_{L^{\infty}(\mathbb{R}^{n+1})}<\infty,$
then $E$ will be removable if and only if the $1/2$-caloric capacity of $E$ is null. One of the main features that characterized the case $s=1/2$ is that the authors could work with an explicit fundamental solution for the operator $\Theta^{1/2}$, namely
@@
$\|P\ast T\|_{L^{\infty}(\mathbb{R}^{n+1})}\leq 1\quad\text{ and }\quad\|P^{\ast}\ast T\|_{L^{\infty}(\mathbb{R}^{n+1})}\leq 1,$
@@
For $E\subset\mathbb{R}^{2}$ compact,
@@
Such result implies, bearing in mind the results of [He], the semi-additivity of $\widetilde{\gamma}_{\Theta^{1/2}}$ in $\mathbb{R}^{2}$ (Theorem 6.8). We also give a result in the general $\mathbb{R}^{n+1}$ setting, also proved in Section 6, where the same type of estimate is established for the typical multidimensional corner-like Cantor set. Sections 2, 3, 4 and 5 are devoted to present the notation and preliminary results necessary to prove the above estimates. In Section 2 we introduce basic terminology and properties, in Section 3 we prove Theorem 3.5, which encapsulates all the different ways to define, equivalently, $\gamma_{\Theta^{1/2},+}$. In Section 4 we use a particular characterization of the latter via a variational approach. With it, we construct a Whitney decomposition of a certain family of compact sets, so that we gain control of their $\widetilde{\gamma}_{\Theta^{1/2}}$ capacity. Finally, in Section 5, we prove Theorem 5.19, a general comparability result between $\widetilde{\gamma}_{\Theta^{1/2}}$ and $\gamma_{\Theta^{1/2}}$ if an additional assumption $\text{{{A}}}_{\boldsymbol{3}}$ is satisfied. In general, the arguments carried out are influenced by those of Tolsa [To3, Ch.5] and Volberg [Vo], where the same type of comparability results are studied for analytic and Lipschitz harmonic capacity respectively.
In our setting, taking into account that $P$ is not an anti-symmetric kernel, such arguments can be applied with some modifications inspired by the arguments presented to prove general $Tb$-theorems such as those found in [NTVo2] and [HyMar]. The fact that $P$ is harmonic outside the hyperplane $\{t=0\}$, an $\pazocal{L}^{n+1}$-null set, and that satisfies being an $n$-dimensional Calderón-Zygmund kernel [MPr, Lemma 2.1] are the essential features used to carry out similar arguments. Finally, Section 7 is devoted to the computation of the $\widetilde{\gamma}_{\Theta^{1/2}}$ capacity of a rectangle $R\subset\mathbb{R}^{2}$, obtaining:
@@
where $\ell_{x}$ and $\ell_{t}$ are the horizontal and vertical side lengths of $R$. The above behavior differs substantially, for example, with that obtained for $\gamma(R)$, the analytic capacity of a rectangle. For the latter one has $\gamma(R)\approx\text{diam}(R)$ (see, for example, [To3, Proposition 1.5]), which is a significantly simpler expression.
Let us also stress one last feature of the above capacities associated to the $1/2$-fractional heat equations: the kernels $P$ and $P^{\ast}$ are both nonnegative. This suggests that, maybe, using classical arguments of potential theory (see [La, Ch.I & II], [Ra, Ch.3] or [Ki]), the comparability between capacities defined through positive measures and distributions should follow by a simpler argument similar to that of [Ve, p.10], provided that one is able to prove the existence of an equilibrium measure. The authors have not been able to deduce its existence for $\gamma_{\Theta^{1/2},+}$, the main obstacle being the incapability of obtaining an equivalent formulation of the latter capacity in terms of an (inverted) infimum over energies. The following simple example already illustrates some problems that arise by proceeding this way: pick $E\subset\mathbb{R}^{2}$ the horizontal line segment $[0,1]\times\{0\}$ and $\mu:=\pazocal{H}^{1}|_{E}$. In [MPr, §6] Mateu & Prat bound explicitly $P\ast\mu$ by $\pi$, obtaining that $\gamma_{\Theta^{1/2},+}(E)\geq\pi^{-1}$. However, by the definition of $P$ and the choice of $\mu$, it is also clear that $P\ast\mu(x,0)=0$, for any $x\in[0,1]$. Therefore, if we were to compute the energy of $\mu$ we would obtain
@@
About the notation used in the sequel: constants appearing in this text can depend on the dimension of the ambient space and possibly on the Calderón-Zygmund (C-Z) constants of $P$, and their value may change at different occurrences. The notation $A\lesssim B$ means that there exists such a constant $C$, so that $A\leq CB$. Moreover, $A\approx B$ is equivalent to $A\lesssim B\lesssim A$. Also, $A\simeq B$ will mean $A=CB$. We will write $B_{r}(x,t)$ to denote the usual Euclidean ball in $\mathbb{R}^{n+1}$ of radius $r>0$ and center $(x,t)$. We also emphasize that the gradient symbol $\nabla$ will refer to $(\nabla_{x},\partial_{t})$, with $x\in\mathbb{R}^{n}$ and $t\in\mathbb{R}$.
@@ 2. Notation and basic definitions
Our ambient space will be $\mathbb{R}^{n+1}$, and a generic point will be denoted as $\overline{x}=(x,t)\in\mathbb{R}^{n+1}$, where $x\in\mathbb{R}^{n}$ will be usually referred to as the spatial variable, and $t\in\mathbb{R}$ the time variable. Most of the results we will present in this text will be valid in a general multidimensional context, i.e. $n$ is any integer bigger or equal than $1$. However, some of them will only hold in the planar setting $(n=1)$, and the authors will make clear when this is the case. Let $\Theta^{1/2}$ be the $1/2$-heat operator,
@@
where $(-\Delta)^{1/2}=(-\Delta_{x})^{1/2}$ is a pseudo-differential operator known as the $1/2$-Laplacian with respect to the spatial variable. It may be defined through its Fourier transform,
$\widehat{(-\Delta)^{1/2}}f(\xi,t)=|\xi|\widehat{f}(\xi,t),$
@@
[…]
$\simeq\int_{\mathbb{R}^{n}}\frac{f(x+y,t)-2f(x,t)+f(x-y,t)}{|y|^{n+1}}\text{d}\pazocal{L}^{n}(y).$
@@
$P(\overline{x})=\frac{t}{\big{(}t^{2}+|x|^{2}\big{)}^{(n+1)/2}}\,\chi_{\{t>0\}}(\overline{x})=\frac{t}{|\overline{x}|^{n+1}}\,\chi_{\{t>0\}}(\overline{x}).$
where $\chi$ is the usual indicator function. Notice that the previous kernel is continuous in $\mathbb{R}^{n+1}\setminus{\{0\}}$ and not differentiable at any point of the hyperplane $\{t=0\}$. Another fundamental function that will frequently appear is $P^{\ast}$,
$P^{\ast}(\overline{x}):=P(-\overline{x})=\frac{-t}{|\overline{x}|^{n+1}}\,\chi_{\{t<0\}}(\overline{x}).$
More generally, for any function $f:\mathbb{R}^{n+1}\to\mathbb{R}$, the notation $f^{\ast}$ will always denote $f^{\ast}(\overline{x}):=f(-\overline{x})$. Hence $f$ being an even function is equivalent to $f=f^{\ast}$. We will also extend this notion for distributions as follows: given $T$ any distribution in $\mathbb{R}^{n+1}$, we define $T^{\ast}$ to be the distribution acting on test functions as
@@
We will say that a distribution is even if $T=T^{\ast}$. It is not hard to check (approximating via test functions and using the associativity of [Ca, Theorem 8.15]) that for any $T$ distribution and $S$ distribution with compact support,
@@
$(-\Delta)^{1/2}$ $P^{\ast}(\overline{x})=C^{\prime}\int_{\mathbb{R}^{n}}\frac{P^{\ast}(x+y,t)-2P^{\ast}(x,t)+P^{\ast}(x-y,t)}{|y|^{n+1}}\text{d}\pazocal{L}^{n}(y)$
$=C^{\prime}\int_{\mathbb{R}^{n}}\frac{P(-x-y,-t)-2P(-x,-t)+P(-x+y,-t)}{|y|^{n+1}}\text{d}\pazocal{L}^{n}(y)=(-\Delta)^{1/2}P(-\overline{x}).$
@@
On the other hand, understanding $\partial_{t}P$ distributionally (that makes sense, since $P$ is a locally integrable function in $\mathbb{R}^{n+1}$ and thus a distribution) we have, for any $\varphi$ test function in $\mathbb{R}^{n+1}$,
@@
Such distributions will be called admissible for $\gamma_{\Theta^{1/2}}(E)$. In [MPr, Lemma 3.4] it is proved that for a distribution $T$ in $\mathbb{R}^{n+1}$ with $\|P\ast T\|_{\infty}\leq 1$, if $\varphi$ is a $\pazocal{C}^{1}$ function supported on $Q\subset\mathbb{R}^{n+1}$ a cube, with $\|\nabla\varphi\|_{\infty}\leq\ell(Q)^{-1}$, then
@@
for some constant $C>0$, that we may assume $C\geq 1$ without loss of generality. If the previous property holds for a distribution $T$, we say that $T$ has $n$-growth (with constant $C$). It is easy to see that the previous definition of growth for a distribution generalizes the usual notion of growth if $T$ coincides with a positive measure. Recall that a positive measure $\mu$ is said to have $n$-growth with constant $C$ if
$\mu\big{(}B(\overline{x},r)\big{)}\leq Cr^{n},\hskip 14.22636pt\text{for all }\,\overline{x}\in\mathbb{R}^{n+1},\,r>0.$
@@
The fact that $C$ depends only on the dimension of the ambient space and the C-Z constants of $P$, motivates the following redefinition of $\gamma_{\Theta^{1/2}}$, that will be the one used throughout the whole text:
Definition 2.1 ($1/2$-caloric capacity).
@@
with the supremum taken over distributions satisfying $\|P\ast T\|_{\infty}\leq 1$ and belonging to $\pazocal{T}(E)$, the set of distributions in $\mathbb{R}^{n+1}$ with $\text{supp}(T)\subset E$ and $n$-growth with constant 1.
@@
$\gamma_{+}(E)=\gamma_{\Theta^{1/2},+}(E):=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma(E),\,\|P\ast\mu\|_{\infty}\leq 1\big{\}},$
where $\Sigma(E)$ is the collection of positive Borel measures supported on $E$ with $n$-growth with constant 1. We also define the following capacity:
$\gamma^{\prime}_{+}(E)=\gamma_{\Theta^{1/2},+}^{\prime}(E):=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma(E),\,\|P\ast\mu\|_{L^{\infty}(\mu)}\leq 1\big{\}}.$
@@
$\overline{\gamma}=\gamma_{\overline{\Theta}^{1/2}},\quad\overline{\gamma}_{+}=\gamma_{\overline{\Theta}^{1/2},+}\quad\text{and}\quad\overline{\gamma}^{\prime}_{+}=\gamma_{\overline{\Theta}^{1/2},+}^{\prime}.$
In some of the results presented in [MPr, §3], one encounters expressions of the form $\langle T,\varphi\rangle$, where $\varphi$ is a compactly supported $\pazocal{C}^{1}$ function, and $T$ is a compactly supported distribution satisfying $\|P\ast T\|_{\infty}\leq 1$. This last estimate and the fact that $(-\Delta_{x})^{1/2}\varphi\in L^{1}(\mathbb{R}^{n+1})$, proved in [MPr, §3], allows to give meaning to $\langle T,\varphi\rangle$ (that a priori may not have sense since $\varphi$ is not a test function) as follows:
$\langle T,\varphi\rangle:=\big{\langle}P\ast T,\overline{\Theta}^{1/2}\varphi\big{\rangle}=\big{\langle}P\ast T,(-\Delta_{x})^{1/2}\varphi+\partial_{t}\varphi\big{\rangle}.$
@@
$\varphi\in\pazocal{C}_{c}(Q)\cap\Big{(}\pazocal{C}_{x}^{m}(Q),\pazocal{C}_{t}^{m}(Q\setminus{\pazocal{N}})\Big{)},$
or simply
$\varphi\in\pazocal{C}^{m}_{c,\pazocal{N}}(Q),$
@@
For $\varphi\in\pazocal{C}^{m}_{c,\pazocal{N}}(Q)$ one can define $\langle T,\varphi\rangle$ as above, assuming $\|P\ast T\|_{\infty}\leq 1$ and $T$ compactly supported. Indeed, observe that we only have to give meaning to $\langle P\ast T,(-\Delta_{x})^{1/2}\varphi\rangle$ and $\langle P\ast T,\partial_{t}\varphi\rangle$. The former term can be defined as in [MPr, §3], where the integrability of $(-\Delta_{x})^{1/2}\varphi$ is proved using only regularity assumptions over the spatial variables. Regarding the term $\langle P\ast T,\partial_{t}\varphi\rangle$, since $P\ast T\in L^{\infty}(\mathbb{R}^{n+1})$, it can be simply defined as
$\langle P\ast T,\partial_{t}\varphi\rangle:=\int_{\text{supp}(\varphi)\setminus{\pazocal{N}}}P\ast T(\overline{x})\partial_{t}\varphi(\overline{x})\mathop{}\!\mathrm{d}\pazocal{L}^{n+1}(\overline{x}),$
since the set $\pazocal{N}$ has null $\pazocal{L}^{n+1}$-measure.
@@
We notice that if for $\varphi\in\pazocal{C}_{c,\pazocal{N}}^{1}(Q)$ it also happens that $\|\nabla\varphi\|_{\infty}\leq\ell(Q)^{-1}$, then [MPr, Corollary 3.3] and [MPr, Lemma 3.5] also hold with exactly the same proofs for such $\varphi$; and therefore also do [MPr, Theorem 3.1] (localization of potentials) and [MPr, Lemma 3.4] ($n$-growth of admissible distributions). Here, condition $\|\nabla\varphi\|_{\infty}\leq\ell(Q)^{-1}$ must be understood as: for every test function $\psi$,
@@
$\widetilde{\gamma}(E)=\widetilde{\gamma}_{\Theta^{1/2}}(E)$ $:=\sup\big{\{}|\langle T,1\rangle|\,:\,T\in\pazocal{T}(E),\,\|P\ast T\|_{\infty}\leq 1,\|P^{\ast}\ast T\|_{\infty}\leq 1\big{\}},$
$\widetilde{\gamma}_{+}(E)=\widetilde{\gamma}_{\Theta^{1/2},+}(E)$ $:=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma(E),\,\|P\ast\mu\|_{\infty}\leq 1,\|P^{\ast}\ast\mu\|_{\infty}\leq 1\big{\}},$
$\widetilde{\gamma}^{\prime}_{+}(E)=\widetilde{\gamma}_{\Theta^{1/2},+}^{\prime}(E)$ $:=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma(E),\,\|P\ast\mu\|_{L^{\infty}(\mu)}\leq 1,\|P^{\ast}\ast\mu\|_{L^{\infty}(\mu)}\leq 1\big{\}}.$
@@
$\gamma_{\text{{sy}},+}(E)$ $:=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma(E),\,\|P_{\text{{sy}}}\ast\mu\|_{\infty}\leq 1\big{\}},$
$\gamma_{\text{{sy}},+}^{\prime}(E)$ $:=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma(E),\,\|P_{\text{{sy}}}\ast\mu\|_{L^{\infty}(\mu)}\leq 1\big{\}},$

### v4

@@ Abstract.
In this paper we study properties of a variant of the $1/2$-caloric capacity, called $1/2$-symmetric caloric capacity. The latter is associated simultaneously with the $1/2$-fractional heat equation and its conjugate. We establish its semi-additivity in $\mathbb{R}^{n+1}$ and, moreover, we compute explicitly the $1/2$-symmetric caloric capacity of rectangles, which illustrates its anisotropic behavior.
@@ 1. Introduction
The study of removable subsets for caloric equations done by Mateu, Prat & Tolsa in [MPrTo] and by Mateu & Prat in [MPr] introduces, naturally, caloric capacities associated to such PDE’s. In the first reference $(1,1/2)$-Lipschitz caloric capacity was introduced, as well as the notion of equivalence between the nullity of this quantity and the removability of compact subsets for the heat equation, i.e. the one associated with the differential operator $\Theta:=(-\Delta_{x})+\partial_{t}$, where $(x,t)\in\mathbb{R}^{n}\times\mathbb{R}$. In [MPr], Mateu & Prat studied the corresponding caloric capacities associated with the fractional heat equation. That is, the equation associated with the pseudo-differential operator $\Theta^{s}:=(-\Delta_{x})^{s}+\partial_{t}$, for $0<s<1$. The authors distinguished the cases $s=1/2$, $1/2<s<1$ and $0<s<1/2$, focusing mainly on the first. The study for $s=1/2$ was quite successful, obtaining a removability result for the $\Theta^{1/2}$-equation, as it was done for the heat equation. For instance, if $f$ is a solution of the $1/2$-heat equation in $\mathbb{R}^{n+1}\setminus{E}$ satisfying $\|f\|_{L^{\infty}(\mathbb{R}^{n+1})}<\infty,$ then $E$ will be removable if and only if the $1/2$-caloric capacity of $E$ is null. One of the main features that characterized the case $s=1/2$ is that the authors could work with an explicit fundamental solution for the operator $\Theta^{1/2}$, namely
@@
[…]
@@
For $E\subset\mathbb{R}^{n+1}$ compact,
@@
Such result implies, bearing in mind the results of [He], the semi-additivity of $\widetilde{\gamma}_{\Theta^{1/2}}$ (Theorem 6.7). Sections 2, 3, 4 and 5 are devoted to present the notation and preliminary results necessary to prove the above estimates. In Section 2 we introduce basic terminology and properties, in Section 3 we prove Theorem 3.5, which encapsulates all the different ways to define, equivalently, $\gamma_{\Theta^{1/2},+}$. In Section 4 we use a particular characterization of the latter via a variational approach. With it, we construct a Whitney decomposition of a certain family of compact sets, so that we gain control of their $\widetilde{\gamma}_{\Theta^{1/2}}$ capacity. Finally, in Section 5, we prove Theorem 5.19, a general comparability result between $\widetilde{\gamma}_{\Theta^{1/2}}$ and $\gamma_{\Theta^{1/2}}$ if an additional assumption $\text{{{A}}}_{\boldsymbol{3}}$ is satisfied, and in Section 6 we remove such assumption and obtain the general result. Our methods are influenced by those of Tolsa [To3, Ch.5] and Volberg [Vo], where the same type of comparability results are studied for analytic and Lipschitz harmonic capacity respectively.
In our setting, since $P$ is not an anti-symmetric kernel, such arguments can be applied with some additional assumptions based on those of general $Tb$-theorems (see [NTVo2] and [HyMar]). The fact that $P$ is harmonic outside the hyperplane $\{t=0\}$, an $\pazocal{L}^{n+1}$-null set, and that satisfies being an $n$-dimensional Calderón-Zygmund kernel [MPr, Lemma 2.1] are the essential features used to carry out similar arguments. Finally, Section 7 is devoted to the computation of the $\widetilde{\gamma}_{\Theta^{1/2}}$ capacity of a rectangle $R\subset\mathbb{R}^{2}$, obtaining:
@@
where $\ell_{x}$ and $\ell_{t}$ are the horizontal and vertical side lengths of $R$. The above behavior differs substantially, for example, with that obtained for $\gamma(R)$, the analytic capacity of a rectangle. For the latter one has $\gamma(R)\approx\text{diam}(R)$ (see [To3, Proposition 1.5]).
Let us also stress one last feature of the above capacities associated to the $1/2$-fractional heat equations: the kernels $P$ and $P^{\ast}$ are both nonnegative. This suggests that, maybe, using classical arguments of potential theory (see [La, Ch.I & II], [Ra, Ch.3] or [Ki]), the comparability between capacities defined through positive measures and distributions should follow by a simpler argument similar to that of [Ve, p.10], provided that one is able to prove the existence of an equilibrium measure. The authors have not been able to deduce its existence for $\gamma_{\Theta^{1/2},+}$, the main obstacle being the incapability of obtaining an equivalent formulation of the latter capacity in terms of an (inverted) infimum over energies. The following simple example already illustrates some problems that arise by proceeding this way: pick $E\subset\mathbb{R}^{2}$ the horizontal line segment $0,1]\times\{0\}$ and $\mu:=\pazocal{H}^{1}|_{E}$. In [MPr, §6] Mateu & Prat bound explicitly $P\ast\mu$ by $\pi$, obtaining that $\gamma_{\Theta^{1/2},+}(E)\geq\pi^{-1}$. However, by the definition of $P$ and the choice of $\mu$, it is also clear that $P\ast\mu(x,0)=0$, for any $x\in[0,1]$. Therefore, if we were to compute the energy of $\mu$ we would obtain
@@
About the notation used in the sequel: We use an upper bar to denote generic points in $\mathbb{R}^{n+1}$, as for example $\overline{x},\overline{y},\overline{z}\ldots$ For a given $\overline{x}=(x,t)$, $x\in\mathbb{R}^{n}$ will be usually referred to as the spatial variable, and $t\in\mathbb{R}$ the time variable.
Constants appearing in this text can depend on the dimension of the ambient space and possibly on the Calderón-Zygmund (CZ) constants of $P$, and their value may change at different occurrences. The notation $A\lesssim B$ means that there exists such a constant $C$, so that $A\leq CB$. Moreover, $A\approx B$ is equivalent to $A\lesssim B\lesssim A$. Also, $A\simeq B$ will mean $A=CB$.
We will write $B_{r}(\overline{x})$ to denote the usual Euclidean ball in $\mathbb{R}^{n+1}$ of radius $r>0$ and center $\overline{x}$. A cube in $\mathbb{R}^{n+1}$ centered at $\overline{x}$ with side length $\ell(Q)$ will be denoted $Q(\overline{x},\ell(Q))$ and usually named $Q,R,S\ldots$ Cubes will always have sides parallel to the coordinate axes. If we write $\lambda Q$ we mean a dilation of $Q$ of factor $\lambda$. That is, $\lambda Q$ will be the cube concentric with $Q$ of side length $\lambda\ell(Q)$ (we follow analogous conventions with balls).
We also emphasize that the gradient symbol $\nabla$ will refer to $(\nabla_{x},\partial_{t})$, with $x\in\mathbb{R}^{n}$ and $t\in\mathbb{R}$.
We will also write $\|\cdot\|_{\infty}:=\|\cdot\|_{L^{\infty}(\mathbb{R}^{n+1})}$.
The symbol $\chi_{A}$, where $A\subset\mathbb{R}^{n+1}$ is Borel measurable, means the indicator function that equals $1$ on $A$, and is 0 otherwise.
For any function $f:\mathbb{R}^{n+1}\to\mathbb{R}$, the symbol $f^{\ast}$ will denote $f^{\ast}(\overline{x}):=f(-\overline{x})$. Hence $f$ being an even function is equivalent to $f=f^{\ast}$.
Fourier transforms of smooth functions $f:\mathbb{R}^{n+1}\to\mathbb{R}$ will be denoted by $\pazocal{F}[f]$ or simply as $\widehat{f}$. They will always be taken with respect to spatial variables. For a fixed $t\in\mathbb{R}$, we put $g_{t}(x):=f(x,t)$ and write $\widehat{f}(\cdot,t):=\widehat{g_{t}}$.
Since the Laplacian operator (fractional or not) will frequently appear in our discussion and will always be taken with respect to spatial variables, we will adopt the notation $\Delta:=\Delta_{x}$.
@@ 2. Notation and basic definitions
Our ambient space will be $\mathbb{R}^{n+1}$ and $\Theta^{1/2}$ will denote the $1/2$-heat operator,
@@
where $(-\Delta)^{1/2}$ is a pseudo-differential operator known as the $1/2$-Laplacian with respect to the spatial variable. It may be defined through its Fourier transform for every fixed $t$,
$\pazocal{F}[{(-\Delta)^{1/2}}f](\xi,t)=|\xi|\pazocal{F}[f](\xi,t),$
@@
$(-\Delta)^{1/2}f(x,t)$ $\simeq\text{p.v.}\int_{\mathbb{R}^{n}}\frac{f(x,t)-f(y,t)}{|x-y|^{n+1}}\text{d}y$
$\simeq\int_{\mathbb{R}^{n}}\frac{f(x+y,t)-2f(x,t)+f(x-y,t)}{|y|^{n+1}}\text{d}y.$
@@
$P(\overline{x})=\frac{t}{|\overline{x}|^{n+1}}\,\chi_{\{t>0\}}(\overline{x}).$
Notice that the previous kernel is locally integrable in $\mathbb{R}^{n+1}$, continuous in $\mathbb{R}^{n+1}\setminus{\{0\}}$ and not differentiable at any point of the hyperplane $\{t=0\}$. Another function that will frequently appear is $P^{\ast}$,
$P^{\ast}(\overline{x}):=P(-\overline{x})=\frac{-t}{|\overline{x}|^{n+1}}\,\chi_{\{t<0\}}(\overline{x}),$
that is the conjugate kernel associated to $P$. For a given distribution $T$, we also define $T^{\ast}$ to be the distribution acting as
@@
It is not hard to check (approximating via test functions and using the associativity of [Ca, Theorem 8.15]) that for any $T$ distribution and $S$ distribution with compact support,
@@
$(-\Delta)^{1/2}$ $P^{\ast}(\overline{x})=C^{\prime}\int_{\mathbb{R}^{n}}\frac{P^{\ast}(x+y,t)-2P^{\ast}(x,t)+P^{\ast}(x-y,t)}{|y|^{n+1}}\text{d}y$
$\hskip-5.69046pt=C^{\prime}\int_{\mathbb{R}^{n}}\frac{P(-x-y,-t)-2P(-x,-t)+P(-x+y,-t)}{|y|^{n+1}}\text{d}y=(-\Delta)^{1/2}P(-\overline{x}).$
@@
On the other hand, we also have for any $\varphi$ test function in $\mathbb{R}^{n+1}$,
@@
Such distributions will be called admissible for $\gamma_{\Theta^{1/2}}(E)$. In [MPr, Lemma 3.4] it is proved that for a distribution $T$ in $\mathbb{R}^{n+1}$ with $\|P\ast T\|_{\infty}\leq 1$, if $\varphi$ is a test function supported on $Q\subset\mathbb{R}^{n+1}$ a cube, with $\|\nabla\varphi\|_{\infty}\leq\ell(Q)^{-1}$, then
@@
for some constant $C>0$, that we may assume $C\geq 1$ without loss of generality.
Definition 2.1.
If the previous property holds for a distribution $T$, we say that $T$ has growth of degree $n$ or simply $n$-growth (with constant $C$). We write $\pazocal{T}_{n}(E)$ those distributions with $n$-growth and constant 1 supported on a compact set $E\subset\mathbb{R}^{n+1}$.
It is easy to see that the previous definition of growth for a distribution generalizes the usual notion of growth if $T$ coincides with a positive measure.
Definition 2.2.
Recall that a positive measure $\mu$ is said to have $n$-growth with constant $C$ if $\mu\big{(}B(\overline{x},r)\big{)}\leq Cr^{n}$, for all $\overline{x}\in\mathbb{R}^{n+1},\,r>0$. We also write $\Sigma_{n}(E)$ the collection measures with $n$-growth and constant 1 supported on a compact set $E\subset\mathbb{R}^{n+1}$.
@@
The fact that $C$ depends only on the dimension of the ambient space and the CZ constants of $P$, motivates the following redefinition of $\gamma_{\Theta^{1/2}}$, that will be the one used throughout the whole text:
Definition 2.3.
@@
the supremum taken over distributions in $\pazocal{T}_{n}(E)$ and satisfying $\|P\ast T\|_{\infty}\leq 1$.
@@
$\gamma_{+}(E)=\gamma_{\Theta^{1/2},+}(E):=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma_{n}(E),\,\|P\ast\mu\|_{\infty}\leq 1\big{\}},$
@@
$\overline{\gamma}=\gamma_{\overline{\Theta}^{1/2}}\quad\text{and}\quad\overline{\gamma}_{+}=\gamma_{\overline{\Theta}^{1/2},+}.$
Moreover, it is usual to extend all of the above definitions to a greater variety of sets. Namely, if $E\subseteq\mathbb{R}^{n+1}$ is any Borel set,
$\gamma(E):=\sup_{\begin{subarray}{c}K\subseteq E\\ K\text{compact}\end{subarray}}\gamma(K),$
and similarly for the rest of capacities.
In some of the results presented in [MPr, §3], one encounters expressions of the form $\langle T,\varphi\rangle$, where $\varphi$ is a compactly supported $\pazocal{C}^{1}$ function, and $T$ is a compactly supported distribution satisfying $\|P\ast T\|_{\infty}\leq 1$. This last estimate and the fact that $(-\Delta)^{1/2}\varphi\in L^{1}(\mathbb{R}^{n+1})$, proved in [MPr, §3], allows to give meaning to $\langle T,\varphi\rangle$ (that a priori may not have sense since $\varphi$ is not a test function) as follows:
$\langle T,\varphi\rangle:=\big{\langle}P\ast T,\overline{\Theta}^{1/2}\varphi\big{\rangle}=\big{\langle}P\ast T,(-\Delta)^{1/2}\varphi+\partial_{t}\varphi\big{\rangle}.$
@@
$\varphi\in\pazocal{C}_{c}(Q)\cap\Big{(}\pazocal{C}_{x}^{m}(Q),\pazocal{C}_{t}^{m}(Q\setminus{\pazocal{N}})\Big{)}\quad\text{or simply}\quad\varphi\in\pazocal{C}^{m}_{c,\pazocal{N}}(Q),$
@@
For $\varphi\in\pazocal{C}^{m}_{c,\pazocal{N}}(Q)$ one can define $\langle T,\varphi\rangle$ as above, assuming $\|P\ast T\|_{\infty}\leq 1$ and $T$ compactly supported. Indeed, observe that we only have to give meaning to $\langle P\ast T,(-\Delta)^{1/2}\varphi\rangle$ and $\langle P\ast T,\partial_{t}\varphi\rangle$. The former term can be defined as in [MPr, §3], where the integrability of $(-\Delta)^{1/2}\varphi$ is proved using only regularity assumptions over the spatial variables. Regarding the term $\langle P\ast T,\partial_{t}\varphi\rangle$, since $P\ast T\in L^{\infty}(\mathbb{R}^{n+1})$, it can be simply defined as
$\langle P\ast T,\partial_{t}\varphi\rangle:=\int_{\text{supp}(\varphi)\setminus{\pazocal{N}}}P\ast T(\overline{x})\partial_{t}\varphi(\overline{x})\mathop{}\!\mathrm{d}\overline{x}.$
@@
We notice that if for $\varphi\in\pazocal{C}_{c,\pazocal{N}}^{1}(Q)$ it also happens that $\|\nabla\varphi\|_{\infty}\leq\ell(Q)^{-1}$, then [MPr, Corollary 3.3] also holds with exactly the same proof for such $\varphi$; and therefore also does [MPr, Theorem 3.1] (localization of potentials) and [MPr, Lemma 3.4] ($n$-growth of admissible distributions). Here, condition $\|\nabla\varphi\|_{\infty}\leq\ell(Q)^{-1}$ must be understood as: for every test function $\psi$,
@@
$\widetilde{\gamma}(E)=\widetilde{\gamma}_{\Theta^{1/2}}(E)$ $:=\sup\big{\{}|\langle T,1\rangle|\,:\,T\in\pazocal{T}_{n}(E),\,\|P\ast T\|_{\infty}\leq 1,\|P^{\ast}\ast T\|_{\infty}\leq 1\big{\}},$
$\widetilde{\gamma}_{+}(E)=\widetilde{\gamma}_{\Theta^{1/2},+}(E)$ $:=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma_{n}(E),\,\|P\ast\mu\|_{\infty}\leq 1,\|P^{\ast}\ast\mu\|_{\infty}\leq 1\big{\}}.$
@@
$\gamma_{\text{{sy}},+}(E)$ $:=\sup\big{\{}\mu(E)\,:\,\mu\in\Sigma_{n}(E),\,\|P_{\text{{sy}}}\ast\mu\|_{\infty}\leq 1\big{\}},$